OpenStudy (anonymous):

solve -3(5-15n)=15(3n-2)

OpenStudy (johnweldon1993):

Use the distributive property...it states x(y+z) = xy + xz Apply that to THIS problem use -3 as x....5 as y...and 15n as z and then use 15 as x...3n as y...and 2 as z

OpenStudy (johnweldon1993):

How'd you do?

OpenStudy (anonymous):

not so good

OpenStudy (anonymous):

hey can you help me

OpenStudy (johnweldon1993):

Alright here's a walkthrough we have -3(5-15n)=15(3n-2) from that distributive property we have (-3*5) (-3 * -15n) = (15*3n) - (15*-2) -15 + 45n = 45n - 30 Now ask yourself...for any number of N you can think of....can n - 30 and n - 15 ever be equal? No, this system have NO solution

OpenStudy (johnweldon1993):

has* not have lol....and heyyyyy whats your question?

OpenStudy (anonymous):

Kevin and Conner tracked their exercise last week. Kevin ran 15 laps around the school track, 10 laps around his neighborhood, and 2 miles on the treadmill. Conner ran 10 laps around the school track and 2 miles on the treadmill. Write and simplify an expression for the total combined distance that Kevin and Conner ran. Show your work. Use t for the distance in laps around their school track and n for the distance in laps around Kevin’s neighborhood.

OpenStudy (anonymous):

smart didnt think of that

OpenStudy (johnweldon1993):

Glad to help :) and heyyyyy I'll reply to your question. go to your question you posted

OpenStudy (anonymous):

ok thanks